# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ 3.3 Conjugacy properties of sue distributions in multivariate censored models
$\begin{split}&{\bm{\xi}}_{\tiny\mbox{nu}}=\begin{bmatrix}{\bm{\xi}}_{\bm{\beta}}+\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})\\ {\bm{\xi}}_{{\bf{y}}_{0}}{+}\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})\end{bmatrix}=:\begin{bmatrix}{\bm{\xi}}_{\tiny\mbox{nu}{\bm{\beta}}}\\ {\bm{\xi}}_{\tiny\mbox{nu}{\bf{y}}_{0}}\end{bmatrix},\quad\bm{\Omega}_{\tiny\mbox{nu}}=\begin{bmatrix}\bm{\Omega}_{{\bm{\beta}}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bm{\beta}}}&\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{0}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}\\ \bm{\Omega}_{{\bf{y}}_{0}{\bm{\beta}}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bm{\beta}}}&\bm{\Omega}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}\end{bmatrix}=:\begin{bmatrix}\bm{\Omega}_{\tiny\mbox{nu}{\bm{\beta}}}&\bm{\Omega}_{\tiny\mbox{nu}{\bm{\beta}}{\bf{y}}_{0}}\\ \bm{\Omega}_{\tiny\mbox{nu}{\bf{y}}_{0}{\bm{\beta}}}&\bm{\Omega}_{\tiny\mbox{nu}{\bf{y}}_{0}}\end{bmatrix},\\ &{\bm{\Delta}}_{\tiny\mbox{nu}}=\begin{bmatrix}\bm{\omega}^{-1}_{\tiny\mbox{nu}\tiny{\bm{\beta}}}(\bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{{\bm{\beta}}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}})\bm{\gamma}^{-1}_{\tiny\mbox{nu}}\\ \bm{\omega}^{-1}_{\tiny\mbox{nu}{\bf{y}}_{0}}(\bm{\omega}_{{\bf{y}}_{0}}{\bm{\Delta}}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}})\bm{\gamma}^{-1}_{\tiny\mbox{nu}}\end{bmatrix}=:\begin{bmatrix}{\bm{\Delta}}_{\tiny\mbox{nu}{\bm{\beta}}}\\ {\bm{\Delta}}_{\tiny\mbox{nu}{\bf{y}}_{0}}\end{bmatrix},\quad\begin{matrix}\phantom{-1234}\bm{\tau}_{\tiny\mbox{nu}}=\bm{\gamma}^{-1}_{\tiny\mbox{nu}}[\bm{\tau}+{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}^{-1}_{{\bf{y}}_{1}}\bm{\omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})],\\ \bar{\bm{\Gamma}}_{\tiny\mbox{nu}}=\bm{\gamma}^{-1}_{\tiny\mbox{nu}}(\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}_{{\bf{y}}_{1}}^{-1}{\bm{\Delta}}_{{\bf{y}}_{1}})\bm{\gamma}^{-1}_{\tiny\mbox{nu}},\end{matrix}\end{split}$
@@
Proposition 3 states a highly–general result that establishes sue conjugacy for a broad class of models whose likelihood factorizes as the product of multivariate elliptical densities and cumulative distribution functions. These likelihoods substantially extend classical tobit representations to multivariate and skewed contexts while covering a broader family of noise terms beyond the Gaussian ones. As clarified in Examples 56, albeit general, such a result allows the recovery of Bayesian formulations of potential interest in practice while ensuring conjugacy under these representations.
Example 5 (Multivariate unified skew–normal (sun) conjugacy).
@@
Classical tobit models consider $y_{i}=\bar{y}_{i}1(\bar{y}_{i}>0)$ with $(\bar{y}_{i}\mid{\bm{\beta}})\sim\mbox{N}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2})$, independently for $i\in\{1,\ldots,n\}$. A natural extension that incorporates skewness within these representations replaces $\mbox{N}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2})$ with ${\mbox{SN}}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2},\alpha)$ [44]. Under this setting, which includes the classical and routinely–implemented tobit formulation when $\alpha=0$, Anceschi et al. [4] have shown that sun priors for ${\bm{\beta}}$, i.e., ${\bm{\beta}}\sim\mbox{SUN}_{p,q}({\bm{\xi}}_{{\bm{\beta}}},\bm{\Omega}_{{\bm{\beta}}},{\bm{\Delta}}_{{\bm{\beta}}},\bm{\tau}_{{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bm{\beta}}})$, yield posterior distributions within the same class. Such a result can be derived as a very special case of Proposition 3 under a Gaussian density generator and suitable constraints on the parameters.
@@
To clarify the above point, assume again the case $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUN}}}_{p+n,q+n}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}})$ with parameters partitioned as in Example 1. Moreover, consider the partitioning $({\bf{y}}^{\top}_{1},{\bf{y}}^{\top}_{0})^{\top}$ defined in Proposition 3 for a generic realization ${\bf{y}}$ from model (25). This construction, combined with the results in Equation (27) and the proof of Proposition 3, implies that $({\bm{\beta}}^{\top},\bar{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUN}}}_{p+n_{1}+n_{0},q+n_{1}+n_{0}}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n_{1}+n_{0}+q+n_{1}+n_{0})}),$ with
@@
$\begin{split}{\bm{\xi}}^{\dagger}=\begin{bmatrix}{\bm{\xi}}_{\bm{\beta}}\\ {\bf{X}}_{1}{\bm{\xi}}_{\bm{\beta}}\\ {\bf{X}}_{0}{\bm{\xi}}_{\bm{\beta}}\end{bmatrix},\quad{\bm{\Delta}}^{\dagger}=\begin{bmatrix}{\bm{\Delta}}_{\bm{\beta}}&\bm{0}&\bm{0}\\ \bm{\omega}_{{\bf{y}}_{1}}^{-1}{\bf{X}}_{1}\bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{\bm{\beta}}&\bm{\omega}_{{\bf{y}}_{1}}^{-1}\sigma\bar{\alpha}{\bf{I}}_{n_{1}}&\bm{0}\\ \ \bm{\omega}_{{\bf{y}}_{0}}^{-1}{\bf{X}}_{0}\bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{\bm{\beta}}&\bm{0}&\bm{\omega}_{{\bf{y}}_{0}}^{-1}\sigma\bar{\alpha}{\bf{I}}_{n_{0}}\end{bmatrix},\quad\bm{\Omega}^{\dagger}=\begin{bmatrix}\bm{\Omega}_{\bm{\beta}}&\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{1}^{\top}&\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{0}^{\top}\\ {\bf{X}}_{1}\bm{\Omega}_{\bm{\beta}}&{\bf{X}}_{1}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{1}^{\top}+\sigma^{2}{\bf{I}}_{n_{1}}&{\bf{X}}_{1}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{0}^{\top}\\ {\bf{X}}_{0}\bm{\Omega}_{\bm{\beta}}&{\bf{X}}_{0}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{1}^{\top}&{\bf{X}}_{0}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{0}^{\top}+\sigma^{2}{\bf{I}}_{n_{0}}\end{bmatrix},\end{split}$
@@
which coincide with the sun prior and skew–normal noise vector for the extension of the tobit model analyzed in the Supplementary Materials of Anceschi et al. [4]. In addition, leveraging Lemma 2 and Remark 2, we have $(\bar{{\bf{y}}}_{1}\mid{\bm{\beta}})\sim\mbox{SUN}_{n_{1},n_{1}}({\bf{X}}_{1}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n_{1}},\bar{\alpha}{\bf{I}}_{n_{1}},\bm{0},{\bf{I}}_{n_{1}})$. Similarly, by the properties of Gaussian density generators, it follows that under the above constraints for the sun parameters, $(\bar{{\bf{y}}}_{0}\perp\bar{{\bf{y}}}_{1}\mid{\bm{\beta}})$. Therefore, $\mathcal{p}(\bar{{\bf{y}}}_{0}\mid\bar{{\bf{y}}}_{1}={\bf{y}}_{1},{\bm{\beta}})=\mathcal{p}(\bar{{\bf{y}}}_{0}\mid{\bm{\beta}})$, and hence, by the same derivations that led to the sun for $(\bar{{\bf{y}}}_{1}\mid{\bm{\beta}})$, we obtain $(\bar{{\bf{y}}}_{0}{\mid}{\bm{\beta}})\sim\mbox{SUN}_{n_{0},n_{0}}({\bf{X}}_{0}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n_{0}},\bar{\alpha}{\bf{I}}_{n_{0}},\bm{0},{\bf{I}}_{n_{0}})$. Combining these results with Proposition 3, Lemma 2 and Remark 2, and recalling the expression for the sun density and cumulative distribution function (see e.g., Arellano-Valle and Azzalini [6]), leads to
@@
which coincides again with the likelihood in Anceschi et al. [4]. As for the models studied in Examples 1 and 3, the above representation also includes several formulations of direct interest in practice. In particular, setting $\alpha=~{}0$ yields classical tobit regression, whereas replacing $\sigma^{2}{\bf{I}}_{n}$ with a full covariance matrix allows to recover multivariate extensions of tobit models, including those based on skewed link functions.
Example 6 concludes our analysis by clarifying that similar, but yet–unexplored, conjugacy properties can be established also when the focus is on models for Student’s $t$ or skew–$t$ censored observations.
Example 6 (Multivariate unified skew–$t$ (sut) conjugacy).
Conjugacy properties for generalizations of tobit models relying on Student’s $t$ or skew–$t$ censored observations are currently lacking. As stated in Corollary 3, these properties can be derived as special cases of Proposition 3 under Student’s $t$ density generators.
@@
The proof of Corollary 3 requires replacing the generic density generators in Proposition 3 with those of the Student’s $t$ and then leveraging the properties of such generators described in Section 2.2.2. ∎
@@
and therefore, under (25), the model underlying a generic observation ${\bf{y}}=(\bar{{\bf{y}}}^{\top}_{1},\bar{{\bf{y}}}^{\top}_{0})^{\top}$ coincides with a multivariate extension of tobit regression having unified skew–$t$ error terms, and Student’s $t$ prior for ${\bm{\beta}}$ uncorrelated with the noise vector ${\bm{\varepsilon}}$. Applying Corollary 3 to such a formulation yields the likelihood
@@
where $\mathcal{p}(\bar{{\bf{y}}}_{1}={\bf{y}}_{1}\mid{\bm{\beta}})$ coincides with the density function, computed at ${\bf{y}}_{1}$, of the ${{\mbox{SUT}}}_{n_{1},q}({\bf{X}}_{1}{\bm{\beta}},\alpha_{{\bm{\beta}}}\bm{\Omega}_{{\bm{\varepsilon}}_{1}},{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}},\bm{0},\bar{\bm{\Gamma}},\nu+p)$ variable, whereas the quantity $\mathbb{P}(\bar{{\bf{y}}}_{0}\leq\bm{0}\mid\bar{{\bf{y}}}_{1}={\bf{y}}_{1},{\bm{\beta}})$ corresponds to the cumulative distribution function, evaluated at $\bm{0}$, of the ${{\mbox{SUT}}}_{n_{0},q}({\bf{X}}_{0}{\bm{\beta}},\alpha_{\bm{\eta}_{-{\bf{y}}_{0}}}\bm{\Omega}_{{\bm{\varepsilon}}_{0}},{\bm{\Delta}}_{{\bm{\varepsilon}}_{0}}\bm{\gamma}^{-1}_{0},\alpha^{-1/2}_{\bm{\eta}_{-{\bf{y}}_{0}}}\bm{\gamma}^{-1}_{0}{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}}^{\top}\bar{\bm{\Omega}}_{{\bm{\varepsilon}}_{1}}^{-1}\bm{\omega}_{{\bm{\varepsilon}}_{1}}^{-1}({\bf{y}}_{1}-{\bf{X}}_{1}{\bm{\beta}}),\bm{\gamma}_{0}^{-1}(\bar{\bm{\Gamma}}-{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}}^{\top}\bar{\bm{\Omega}}_{{\bm{\varepsilon}}_{1}}^{-1}{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}})\bm{\gamma}_{0}^{-1},\nu+p+n_{1})$, with $\bm{\gamma}_{0}$ defined as $\bm{\gamma}_{0}=\mbox{diag}(\bar{\bm{\Gamma}}-{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}}^{\top}\bar{\bm{\Omega}}_{{\bm{\varepsilon}}_{1}}^{-1}{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}})^{1/2}$. This result clarifies that classical multivariate tobit representations admit extensions to suitable skew–$t$ formulations while preserving conjugacy. Imposing additional constraints within such a formulation further highlights the practical potential of our contribution. For example, setting ${\bm{\Delta}}_{{\bm{\varepsilon}}}=\bm{0}$ in the above formulation, and recalling again Examples 2 and 4, yields

### v3

@@ 3.3 Conjugacy properties of SUE distributions in multivariate censored models
$\begin{split}&{\bm{\xi}}_{\tiny\mbox{nu}}{=}\ \begin{bmatrix}{\bm{\xi}}_{\bm{\beta}}+\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})\\ {\bm{\xi}}_{{\bf{y}}_{0}}{+}\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})\end{bmatrix}=:\begin{bmatrix}{\bm{\xi}}_{\tiny\mbox{nu}{\bm{\beta}}}\\ {\bm{\xi}}_{\tiny\mbox{nu}{\bf{y}}_{0}}\end{bmatrix},\ \ \bm{\Omega}_{\tiny\mbox{nu}}{=}\ \begin{bmatrix}\bm{\Omega}_{{\bm{\beta}}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bm{\beta}}}&\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{0}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}\\ \bm{\Omega}_{{\bf{y}}_{0}{\bm{\beta}}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bm{\beta}}}&\bm{\Omega}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\Omega}_{{\bf{y}}_{1}{\bf{y}}_{0}}\end{bmatrix}=:\begin{bmatrix}\bm{\Omega}_{\tiny\mbox{nu}{\bm{\beta}}}&\bm{\Omega}_{\tiny\mbox{nu}{\bm{\beta}}{\bf{y}}_{0}}\\ \bm{\Omega}_{\tiny\mbox{nu}{\bf{y}}_{0}{\bm{\beta}}}&\bm{\Omega}_{\tiny\mbox{nu}{\bf{y}}_{0}}\end{bmatrix},\\ &{\bm{\Delta}}_{\tiny\mbox{nu}}{=}\ \begin{bmatrix}\bm{\omega}^{-1}_{\tiny\mbox{nu}\tiny{\bm{\beta}}}(\bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{{\bm{\beta}}}-\bm{\Omega}_{{\bm{\beta}}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}})\bm{\gamma}^{-1}_{\tiny\mbox{nu}}\\ \bm{\omega}^{-1}_{\tiny\mbox{nu}{\bf{y}}_{0}}(\bm{\omega}_{{\bf{y}}_{0}}{\bm{\Delta}}_{{\bf{y}}_{0}}-\bm{\Omega}_{{\bf{y}}_{0}{\bf{y}}_{1}}\bm{\Omega}^{-1}_{{\bf{y}}_{1}}\bm{\omega}_{{\bf{y}}_{1}}{\bm{\Delta}}_{{\bf{y}}_{1}})\bm{\gamma}^{-1}_{\tiny\mbox{nu}}\end{bmatrix}=:\begin{bmatrix}{\bm{\Delta}}_{\tiny\mbox{nu}{\bm{\beta}}}\\ {\bm{\Delta}}_{\tiny\mbox{nu}{\bf{y}}_{0}}\end{bmatrix},\quad\begin{matrix}\qquad\bm{\tau}_{\tiny\mbox{nu}}=\bm{\gamma}^{-1}_{\tiny\mbox{nu}}[\bm{\tau}+{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}^{-1}_{{\bf{y}}_{1}}\bm{\omega}^{-1}_{{\bf{y}}_{1}}({\bf{y}}_{1}-{\bm{\xi}}_{{\bf{y}}_{1}})],\\ \bar{\bm{\Gamma}}_{\tiny\mbox{nu}}=\bm{\gamma}^{-1}_{\tiny\mbox{nu}}(\bar{\bm{\Gamma}}-{\bm{\Delta}}^{\top}_{{\bf{y}}_{1}}\bar{\bm{\Omega}}_{{\bf{y}}_{1}}^{-1}{\bm{\Delta}}_{{\bf{y}}_{1}})\bm{\gamma}^{-1}_{\tiny\mbox{nu}},\end{matrix}\end{split}$
@@
Proposition 3 states a general result that establishes SUE conjugacy for a broad class of models whose likelihood factorizes as the product of multivariate elliptical densities and cumulative distribution functions. These likelihoods substantially extend classical tobit representations to multivariate and skewed contexts while covering a broader family of noise terms beyond the Gaussian ones. As clarified in Examples 56, albeit general, such a result allows the recovery of Bayesian formulations of potential interest in practice while ensuring conjugacy under these representations.
Example 5 (Multivariate unified skew-normal (SUN) conjugacy).
@@
Classical tobit models consider $y_{i}=\bar{y}_{i}1(\bar{y}_{i}>0)$ with $(\bar{y}_{i}\mid{\bm{\beta}})\sim\mbox{N}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2})$, independently for $i\in\{1,\ldots,n\}$. A natural extension that incorporates skewness within these representations replaces $\mbox{N}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2})$ with ${\mbox{SN}}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2},\alpha)$ [44]. Under this setting, which includes the classical and routinely-implemented tobit formulation when $\alpha=0$, Anceschi et al. [4] have shown that SUN priors for ${\bm{\beta}}$, i.e., ${\bm{\beta}}\sim\mbox{SUN}_{p,q}({\bm{\xi}}_{{\bm{\beta}}},\bm{\Omega}_{{\bm{\beta}}},{\bm{\Delta}}_{{\bm{\beta}}},\bm{\tau}_{{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bm{\beta}}})$, yield posterior distributions within the same class. Such a result can be derived as a very special case of Proposition 3 under a Gaussian density generator and suitable constraints on the parameters.
@@
To clarify the above point, assume again the case $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUN}}}_{p+n,q+n}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}})$ with parameters partitioned as in Example 1. Moreover, consider the partitioning $({\bf{y}}^{\top}_{1},{\bf{y}}^{\top}_{0})^{\top}$ defined in Proposition 3 for a generic realization ${\bf{y}}$ from model (25). This construction, combined with the results in (27) and the proof of Proposition 3, implies that $({\bm{\beta}}^{\top},\bar{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUN}}}_{p+n_{1}+n_{0},q+n_{1}+n_{0}}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}}),$ with
@@
$\begin{split}{\bm{\xi}}^{\dagger}=\begin{bmatrix}{\bm{\xi}}_{\bm{\beta}}\\ {\bf{X}}_{1}{\bm{\xi}}_{\bm{\beta}}\\ {\bf{X}}_{0}{\bm{\xi}}_{\bm{\beta}}\end{bmatrix},\ \ {\bm{\Delta}}^{\dagger}=\begin{bmatrix}{\bm{\Delta}}_{\bm{\beta}}&\bm{0}&\bm{0}\\ \bm{\omega}_{{\bf{y}}_{1}}^{-1}{\bf{X}}_{1}\bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{\bm{\beta}}&\bm{\omega}_{{\bf{y}}_{1}}^{-1}\sigma\bar{\alpha}{\bf{I}}_{n_{1}}&\bm{0}\\ \ \bm{\omega}_{{\bf{y}}_{0}}^{-1}{\bf{X}}_{0}\bm{\omega}_{\bm{\beta}}{\bm{\Delta}}_{\bm{\beta}}&\bm{0}&\bm{\omega}_{{\bf{y}}_{0}}^{-1}\sigma\bar{\alpha}{\bf{I}}_{n_{0}}\end{bmatrix},\ \ \bm{\Omega}^{\dagger}=\begin{bmatrix}\bm{\Omega}_{\bm{\beta}}&\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{1}^{\top}&\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{0}^{\top}\\ {\bf{X}}_{1}\bm{\Omega}_{\bm{\beta}}&{\bf{X}}_{1}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{1}^{\top}+\sigma^{2}{\bf{I}}_{n_{1}}&{\bf{X}}_{1}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{0}^{\top}\\ {\bf{X}}_{0}\bm{\Omega}_{\bm{\beta}}&{\bf{X}}_{0}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{1}^{\top}&{\bf{X}}_{0}\bm{\Omega}_{\bm{\beta}}{\bf{X}}_{0}^{\top}+\sigma^{2}{\bf{I}}_{n_{0}}\end{bmatrix},\end{split}$
@@
that coincide with the SUN prior and skew-normal noise vector for the extension of the tobit model analyzed in the supplementary materials of Anceschi et al. [4]. In addition, leveraging Lemma 2 and Remark 2, we have $(\bar{{\bf{y}}}_{1}\mid{\bm{\beta}})\sim\mbox{SUN}_{n_{1},n_{1}}({\bf{X}}_{1}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n_{1}},\bar{\alpha}{\bf{I}}_{n_{1}},\bm{0},{\bf{I}}_{n_{1}})$. Similarly, by the properties of Gaussian density generators, it follows that under the above constraints for the SUN parameters, $(\bar{{\bf{y}}}_{0}\perp\bar{{\bf{y}}}_{1}\mid{\bm{\beta}})$. Therefore, $\mathcal{p}(\bar{{\bf{y}}}_{0}\mid\bar{{\bf{y}}}_{1}={\bf{y}}_{1},{\bm{\beta}})=\mathcal{p}(\bar{{\bf{y}}}_{0}\mid{\bm{\beta}})$, and hence, by the same derivations that led to the SUN for $(\bar{{\bf{y}}}_{1}\mid{\bm{\beta}})$, we obtain $(\bar{{\bf{y}}}_{0}{\mid}{\bm{\beta}})\sim\mbox{SUN}_{n_{0},n_{0}}({\bf{X}}_{0}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n_{0}},\bar{\alpha}{\bf{I}}_{n_{0}},\bm{0},{\bf{I}}_{n_{0}})$. Combining these results with Proposition 3, Lemma 2 and Remark 2, and recalling the expression for the SUN density and cumulative distribution function (see e.g., Arellano-Valle and Azzalini [6]), leads to
@@
which coincides again with the likelihood in Anceschi et al. [4]. As for the models explored in Examples 1 and 3, the above representation also includes several formulations of direct interest in practice. In particular, setting $\alpha=0$ yields classical tobit regression, whereas replacing $\sigma^{2}{\bf{I}}_{n}$ with a full covariance matrix allows to recover multivariate extensions of tobit models, including those based on skewed link functions.
Example 6 concludes our analysis by clarifying that similar, but yet-unexplored, conjugacy properties can be established also when the focus is on models for Student’s $t$ or skew-$t$ censored observations.
Example 6 (Multivariate unified skew-$t$ (SUT) conjugacy).
Conjugacy properties for generalizations of tobit models relying on Student’s $t$ or skew-$t$ censored observations are currently lacking. As stated within Corollary 3, these properties can be derived as special cases of Proposition 3 under Student’s $t$ density generators.
@@
The proof of Corollary 3 requires replacing the generic density generators in Proposition 3 with those of the Student’s $t$, and then leveraging the properties of such generators described in Section 2.2.2. ∎
@@
and therefore, under (25), the model underlying a generic observation ${\bf{y}}=(\bar{{\bf{y}}}^{\top}_{1},\bar{{\bf{y}}}^{\top}_{0})^{\top}$ coincides with a multivariate extension of tobit regression having unified skew-$t$ error terms, and Student’s $t$ prior for ${\bm{\beta}}$ uncorrelated with the noise vector ${\bm{\varepsilon}}$. Applying Corollary 3 to such a formulation yields the likelihood
@@
where $\mathcal{p}(\bar{{\bf{y}}}_{1}={\bf{y}}_{1}\mid{\bm{\beta}})$ coincides with the density function, computed at ${\bf{y}}_{1}$, of the ${{\mbox{SUT}}}_{n_{1},q}({\bf{X}}_{1}{\bm{\beta}},\alpha_{{\bm{\beta}}}\bm{\Omega}_{{\bm{\varepsilon}}_{1}},{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}},\bm{0},\bar{\bm{\Gamma}},\nu+p)$ variable, whereas the quantity $\mathbb{P}(\bar{{\bf{y}}}_{0}\leq\bm{0}\mid\bar{{\bf{y}}}_{1}={\bf{y}}_{1},{\bm{\beta}})$ corresponds to the cumulative distribution function, evaluated at $\bm{0}$, of the ${{\mbox{SUT}}}_{n_{0},q}({\bf{X}}_{0}{\bm{\beta}},\alpha_{\bm{\eta}_{-{\bf{y}}_{0}}}\bm{\Omega}_{{\bm{\varepsilon}}_{0}},{\bm{\Delta}}_{{\bm{\varepsilon}}_{0}}\bm{\gamma}^{-1}_{0},\alpha^{-1/2}_{\bm{\eta}_{-{\bf{y}}_{0}}}\bm{\gamma}^{-1}_{0}{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}}^{\top}\bar{\bm{\Omega}}_{{\bm{\varepsilon}}_{1}}^{-1}\bm{\omega}_{{\bm{\varepsilon}}_{1}}^{-1}({\bf{y}}_{1}-{\bf{X}}_{1}{\bm{\beta}}),\bm{\gamma}_{0}^{-1}(\bar{\bm{\Gamma}}-{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}}^{\top}\bar{\bm{\Omega}}_{{\bm{\varepsilon}}_{1}}^{-1}{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}})\bm{\gamma}_{0}^{-1},\nu+p+n_{1})$, with $\bm{\gamma}_{0}$ defined as $\bm{\gamma}_{0}=\mbox{diag}(\bar{\bm{\Gamma}}-{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}}^{\top}\bar{\bm{\Omega}}_{{\bm{\varepsilon}}_{1}}^{-1}{\bm{\Delta}}_{{\bm{\varepsilon}}_{1}})^{1/2}$. This result clarifies that classical multivariate tobit representations admit extensions to suitable skew-$t$ formulations while preserving conjugacy. Imposing additional constraints within such a formulation further highlights the practical potential of our contribution. For example, setting ${\bm{\Delta}}_{{\bm{\varepsilon}}}=\bm{0}$ in the above formulation, and recalling again Examples 2 and 4, yields